OpenStudy (joshoyen):

use the information provided to write the standard form equation of each parabola. 1. y = 2(x + 7)^2 - 6

OpenStudy (astrophysics):

In this one we're going backwards, what is vertex in this?

OpenStudy (joshoyen):

is it this formula -b/2a ?

OpenStudy (astrophysics):

No remember what i showed you on the last one, vertex form, what is the vertex just by looking at it (h,k)

OpenStudy (joshoyen):

ohhh

OpenStudy (joshoyen):

(-7, -6)

OpenStudy (astrophysics):

Right!

OpenStudy (astrophysics):

Ok was checking if you understood it, you just need to foil to put it in standard form y =ax^2+bx+c

OpenStudy (astrophysics):

\[y=2(x+7)^2-6 \] \[(x+7)^2 = (x+7)(x+7)\]

OpenStudy (joshoyen):

@Astrophysics wait i'm confused, what did you do there?

OpenStudy (astrophysics):

You just have to foil out the (x+7)^2 the exponent 2 means there's two of them

OpenStudy (astrophysics):

 multiply...